# When Does Learning Renormalize? Sufficient Conditions for Power–Law Spectral Dynamics

### v4

@@ 1 Introduction
The remainder of the paper is organized as follows. Section 2 reviews the GRSD framework and the formulation of shell dynamics. Section 3 states the sufficient conditions for renormalizable shell dynamics and presents the main theorem. Section 5 interprets these conditions and discusses their relationship to common deep learning design choices. Finally, Section 8 discusses limitations, non–universal regimes, and directions for future work.
@@ 3.2 Statement of sufficient conditions
Let $s=\log\lambda$ and let $\{\mathcal{B}_{i}\}$ denote logarithmic spectral bins of fixed width $h$ in $s$. On an intermediate spectral window, the renormalized shell coupling statistics entering the GRSD closure admit an asymptotically translation-invariant form in $s$: there exists a kernel $K_{h}(\Delta)$, $\Delta=(j-i)h$, such that
where $\mathrm{err}(n,h,L)\to 0$ in the joint limit of large width, small bin size, and sufficient depth.
Condition 4 requires the absence of any intrinsic absolute scale in the renormalized spectral couplings beyond relative separations in logarithmic scale. It forces that $\dot{s}=\frac{\dot{\lambda}}{\lambda}$ is invariant to the absolute scale of $\lambda$. Unlike Conditions 13, this condition does not follow from generic stability or locality considerations and constitutes a genuinely nontrivial structural requirement.
@@ 7.1 Experimental setup
The residual model is a standard ResNet-18. The plain counterpart (PlainNet-18) is constructed by removing all residual skip connections while keeping convolutional blocks, batch normalization, and parameter counts identical. Both models share the same initialization and trained checkpoints.
@@ 7.4 Residual learning induces Toeplitz spectral structure
7.5 Controlled inter-shell coupling and shell coarse-graining
To further dissect the observed Toeplitz structure, we separate contributions from intra-shell (on-diagonal) and inter-shell (off-diagonal) couplings.
Figure 2 shows that inter-shell couplings are suppressed by several orders of magnitude toward the spectral center, while intra-shell mixing remains comparatively large and stable. This clear separation indicates that cross-shell interactions act as a controlled perturbation, consistent with the presence of effective spectral gaps between shells.
These results justify the shell-based coarse-graining used throughout our theoretical analysis: rapid intra-shell equilibration dominates local dynamics, while inter-shell transport governs slower, structured spectral flow.
Figure 2: Comparison between off-diagonal (inter-shell) and on-diagonal (intra-shell) Toeplitz residuals as a function of spectral window position. Inter-shell couplings are strongly suppressed toward the spectral center, while intra-shell mixing remains comparatively large and stable.
@@ A.2 Proof of Theorem 1
Condition 4 rules out that, on an intermediate spectral window, shell couplings are translation–invariant in $s=\log\lambda$. Let $s=\log\lambda$. Concretely, there exists a kernel $K_{h}$ such that
$\widehat{K}_{ij}=K_{h}\!\big((j-i)h\big)+\mathrm{err}(n,h,L),\qquad s_{i}=ih,$
with
$\mathrm{err}(n,h,L)\to 0$
in the joint limit of large width $n$, small log–bin size $h$, and sufficient depth $L$.

### v5

@@ Abstract
Beyond the theoretical analysis, we provide direct empirical evidence that the key structural requirement—log–shift invariance of renormalized shell couplings—is approximately realized in modern residual architectures, and is significantly degraded in non–residual counterparts. These experiments validate the non–vacuous nature of the sufficient conditions and clarify the architectural mechanisms that promote renormalizable spectral transport.
@@ 1 Introduction
A central difficulty in this program is that some of the required conditions—most notably log–shift invariance of renormalized shell couplings—are genuinely structural and do not follow from generic stability or locality considerations alone. Such conditions would be of limited interest if they were never even approximately realized in practical learning systems. For this reason, an essential part of our analysis is to identify concrete architectural mechanisms that promote these properties, and to empirically test whether they are approximately satisfied in realistic training settings.
@@
At first sight, the sufficient conditions identified in this work may appear strong or even restrictive. Indeed, they impose nontrivial requirements on architectural structure, initialization geometry, and training stability. However, it is precisely these requirements that align strikingly well with several core design principles of modern deep learning systems.
Contemporary architectures are overwhelmingly engineered to avoid uncontrolled recurrence or long–range instantaneous coupling, favoring feedforward or residual structures with bounded gradient propagation. Training pipelines are carefully designed to ensure stability and controllability, with explicit emphasis on preventing gradient explosion or catastrophic spectral reorganization. Likewise, random or weakly correlated initialization schemes are ubiquitous, reflecting an implicit preference for functional incoherence at the start of training. Viewed through the GRSD lens, these widely adopted engineering choices correspond closely to Conditions 13, which enforce locality, stability, and statistical regularity of the induced operator dynamics.
From this perspective, the emergence of neural scaling laws should not be interpreted as an intrinsic or inevitable property of deep learning models. Rather, it reflects the fact that engineering requirements of trainability, stability, and scalability progressively constrain learning systems toward a dynamical regime with no intrinsic spectral scale and a renormalizable coarse–grained description. Once learning dynamics enters this approximately scale–free and renormalizable region, power–law behavior is no longer a matter of modeling choice or empirical coincidence, but instead arises as a rigidity consequence of the underlying dynamics.
In addition to the theoretical analysis, we empirically probe the spectral transport operators induced during training and directly measure the resulting coarse–grained shell couplings. These experiments demonstrate that residual architectures approximately realize the log–shift invariance required by the theory, while structurally similar non–residual architectures do not. This contrast provides direct evidence that the sufficient conditions identified here capture meaningful architectural distinctions, rather than abstract or vacuous assumptions.
@@
The remainder of the paper is organized as follows. Section 2 reviews the GRSD framework and the formulation of shell dynamics. Section 3 states the sufficient conditions for renormalizable shell dynamics and presents the main theorem. Section 4 provides a concrete structural mechanism—based on residual learning—by which the most nontrivial condition can be realized. Section 5 interprets these conditions and discusses their relationship to common deep learning design choices. Section 7 presents empirical validation of the structural assumptions underlying the theory. Finally, Section 8 discusses limitations, non–universal regimes, and directions for future work.
@@ 3.2 Statement of sufficient conditions
Let $s=\log\lambda$, and let $\{B_{i}\}$ denote logarithmic spectral shells of fixed width $h$ in $s$, centered at $s_{i}=ih$. Let $\widehat{\Omega}_{ij}$ denote the shell-level renormalized coupling, obtained by coarse-graining the mode-level operator $\dot{M}$ over shells $B_{i}$ and $B_{j}$.
On an intermediate spectral window, the shell-level statistics satisfy Log–shift invariance, i.e. $\widehat{\Omega}_{ij}$ follows scale-free structure:
$\widehat{\Omega}_{ij}(t)=\begin{cases}\mathcal{K}_{h}\!\left((j-i)h\right)+\mathrm{err}_{ij},&i\neq j,\\[4.30554pt] c\lambda_{i}+\mathrm{err}_{ii},&i=j,\end{cases}$
where $\mathcal{K}_{h}$ depends only on the relative log-scale separation, and the residual terms $\mathrm{err}_{ij}$ vanish in the joint limit of large model size and fine shell resolution.
Condition 4 requires the absence of any intrinsic absolute scale in the renormalized spectral couplings beyond relative separations in logarithmic scale. It forces that $\dot{s}=\frac{\dot{\lambda}}{\lambda}$ is invariant to the absolute scale of $\lambda$ (detailed proof provided in A.2). Unlike Conditions 13, this condition does not follow from generic stability or locality considerations and constitutes a genuinely nontrivial structural requirement.
@@ 7.1 Experimental setup
The residual model is a standard ResNet-18 He et al. (2016). The plain counterpart (PlainNet-18) is constructed by removing all residual skip connections while keeping convolutional blocks, batch normalization, and parameter counts identical. Both models share the same trained parameter checkpoints.
@@ 7.3 Shell binning and Toeplitz diagnostics
7.4 Diagonal scaling of spectral couplings
(a) PlainNet
(b) ResNet
Figure 1: Log–log plots of diagonal spectral couplings $\Omega_{ii}$ versus eigenvalues $\lambda_{i}$ for PlainNet (left) and ResNet (right). Each point corresponds to a log-shell averaged diagonal entry. Solid lines denote linear least-squares fits in logarithmic coordinates. ResNet exhibits a scaling exponent closer to unity and significantly reduced scatter compared to PlainNet, consistent with the scale-free diagonal structure predicted by Condition 4.
We empirically examine the diagonal components of the renormalized spectral coupling matrix, $\Omega_{ii}$, to test the scale-free diagonal structure postulated in Condition 4. Recall that Condition 4 predicts that, on an intermediate spectral window, the diagonal terms satisfy a linear scaling relation
$\Omega_{ii}\propto\lambda_{i},$ (6)
corresponding to scale-invariant drift in logarithmic spectral coordinates.
Figure 1 shows log–log scatter plots of $\Omega_{ii}$ versus $\lambda_{i}$ for both architectures, together with linear least-squares fits in logarithmic coordinates. In both cases, the data exhibit a clear power-law relation over a broad intermediate spectral range, consistent with the scale-free diagonal ansatz of Condition 4.
Quantitatively, the ResNet model displays an almost perfectly linear scaling with fitted slope close to unity and an excellent coefficient of determination ($R^{2}\approx 0.99$). In contrast, while PlainNet also shows approximate power-law behavior, its fitted slope deviates more noticeably from unity and exhibits increased scatter, reflected in a lower $R^{2}$ value. This indicates that the diagonal scaling relation is significantly more stable and coherent in the presence of residual connections.
Interpretation.
The near-linear relation $\Omega_{ii}\sim\lambda_{i}$ observed in ResNet supports the assumption that diagonal spectral couplings introduce no intrinsic absolute scale beyond that already encoded in $\lambda_{i}$. This behavior is a necessary ingredient for scale-invariant log-spectral drift and underpins the emergence of renormalizable spectral dynamics. The degradation of this relation in PlainNet suggests that residual connections play a structural role in stabilizing diagonal scaling, thereby promoting the conditions required for log-shift invariance and controlled spectral transport.
Overall, these results provide direct empirical support for the diagonal component of Condition 4 and highlight the architectural dependence of scale-free spectral structure.
@@ A.2 Proof of Theorem 1
Condition 4 rules out that, on an intermediate spectral window, shell couplings are translation–invariant in $s=\log\lambda$. In other words, it forces that $\dot{s}=\frac{\dot{\lambda}}{\lambda}$ is invariant to the absolute scale of $\lambda$.
Scale-invariance of log-spectral drift induced by Condition 4.
Recall that under Condition 4, the shell-level coupling statistics admit a log-shift invariant form on an intermediate spectral window,
$\hat{\Omega}_{ij}(t)=\begin{cases}K_{h}\big((j-i)h\big),&i\neq j,\\[4.0pt] c\,\lambda_{i},&i=j,\end{cases}$ (35)
where $h$ denotes the log-shell width, $\lambda_{i}=\lambda_{0}e^{-ih}$, and $K_{h}$ depends only on the relative log-scale separation.
The evolution of individual eigenvalues takes the form
$\dot{\lambda}_{i}=\sum_{j\neq i}\frac{\hat{\Omega}_{ij}}{\lambda_{i}-\lambda_{j}}+\hat{\Omega}_{ii}.$ (36)
Introducing the log-spectral coordinate $s_{i}:=\log\lambda_{i}$, we obtain
$\dot{s}_{i}=\frac{\dot{\lambda}_{i}}{\lambda_{i}}=\sum_{j\neq i}\frac{\hat{\Omega}_{ij}}{\lambda_{i}(\lambda_{i}-\lambda_{j})}+\frac{\hat{\Omega}_{ii}}{\lambda_{i}}.$ (37)
The diagonal contribution simplifies immediately as
$\frac{\hat{\Omega}_{ii}}{\lambda_{i}}=c,$ (38)
which is independent of the absolute scale of $\lambda_{i}$.
For the off-diagonal terms, writing $\lambda_{i}-\lambda_{j}=\lambda_{i}\big(1-e^{-(s_{j}-s_{i})}\big)$ and using the log-shift invariant structure $\hat{\Omega}_{ij}=K_{h}(s_{j}-s_{i}),$ we find that each summand depends only on the relative separation $s_{j}-s_{i}$. Upon coarse-graining over shells and passing to the continuum limit, the sum reduces to an integral of the form
$\dot{s}(s,t)=\int\mathrm{d}s^{\prime}\,\frac{K(s^{\prime}-s)}{1-e^{-(s^{\prime}-s)}}+c,$ (39)
where $K$ is a scale-free kernel inherited from $K_{h}$.
Equation (39) depends exclusively on relative log-scale separations and contains no reference to the absolute value of $s$ or $\lambda$. Consequently,
$\dot{s}(s+\Delta,t)=\dot{s}(s,t),\qquad\forall\,\Delta\in\mathbb{R},$ (40)
demonstrating that the log-eigenvalue drift is invariant under global spectral rescalings $\lambda\mapsto e^{\Delta}\lambda$. This establishes explicitly that Condition 4 enforces the absence of any intrinsic absolute spectral scale in the renormalized spectral dynamics.
@@ A.3.5 Log–bin averaging and kernel emergence
On-Diagonal shell statistics
The diagonal structure in Condition 4 can be obtained from the same bulk self-mixing mechanism used in the off-diagonal analysis. Fix an intermediate spectral window and condition on a mode $u$ with $\lambda=\langle u,Mu\rangle$ lying in a log-shell $B_{i}$. Beyond a boundary depth, residual blocks induce fast mixing on the projective space of directions, so that the conditional distribution of mode directions $u\mid(\lambda\in B_{i})$ is asymptotically isotropic (up to a small mixing error).
Consequently, for any layerwise diagonal quadratic observable of the form $\Omega^{(\ell)}_{uu}=\langle u,\dot{M}^{(\ell)}u\rangle$, isotropy implies that its conditional expectation retains only the Rayleigh component:
$\mathbb{E}\!\left[\Omega^{(\ell)}_{uu}\,\middle|\,\lambda\in B_{i}\right]=\alpha_{\ell}(t)\,\mathbb{E}\!\left[\langle u,Mu\rangle\,\middle|\,\lambda\in B_{i}\right]\;+\;O(\eta)\,\mathbb{E}\!\left[\lambda\,\middle|\,\lambda\in B_{i}\right],$
for some scalar $\alpha_{\ell}(t)$ independent of the absolute shell index $i$ on the intermediate window, and where the $O(\eta)$ term captures the finite-depth imperfect mixing (the same $O(\eta)$ level as in the bulk part of the Toeplitz argument). This reduction to the Rayleigh component is standard for isotropic quadratic observables and follows from classical results on quadratic forms under rotational invariance [22, 2, 29].
Summing over layers and applying the same boundary–bulk decomposition as in depth dilution, the boundary contribution is suppressed by $\ell^{*}/L$, while the bulk average preserves the shell-independent coefficient. Hence the full diagonal coupling satisfies, on the intermediate window,
$\mathbb{E}\!\left[\Omega_{uu}\,\middle|\,\lambda\in B_{i}\right]=\big(c(t)+O(\eta)+O(\ell^{*}/L)\big)\,\mathbb{E}\!\left[\lambda\,\middle|\,\lambda\in B_{i}\right],$
i.e. $\Omega_{ii}=c(t)\lambda_{i}+\mathrm{err}_{ii}$ after shell averaging, which is the on-diagonal part of Condition 4. Standard self-averaging over depth and within-shell averaging further reduce the residual fluctuations around this mean.